# Multi-objective Combinatorial Methodology for Nuclear Reactor Site Assessment: A Case Study for the United States

### v1

@@ Acknowledgment
This work is sponsored by the Department of Energy Office of Nuclear Energy’s Distinguished Early Career Program (Award number: DE-NE0009424), which is administered by the Nuclear Energy University Program (NEUP). This research also made use of Idaho National Laboratory computing resources, which are supported by the Office of Nuclear Energy of the U.S. Department of Energy under Contract No. DE-AC07-05ID14517.
@@
Appendix A Details of the Top Coal and Brownfield Sites
@@
Registry ID     110015334440 110070068631 C56068 110070068960 C2712 C10343
Longitude     -122.1032 -122.0370 -87.8336 -120.9098 -79.0731 -76.4530
Latitude     37.6677 38.7013 42.8492 40.0949 36.4833 40.8112
County & State     Almeda, CA Yolo, CA Milwaukee, WI Plumas, CA Person, NC Northumberland, PA
Siting Metric     1.0000 0.8506 0.8211 0.7703 0.7614 0.7501
Merged Dataset Rank     1 2 3 4 5 6
Number of Nuclear Restrictions in the State     1 1 0 1 0 0
State Electricity Price     24.2231 24.2231 26.0478 24.2231 14.0500 14.0500
State Net Electricity Imports     75504 75504 12109 75504 14875 -70484
State Nuclear Inclusive Policy     1 1 0 1 1 0
Population Sentiment of Nuclear Energy     0.4019 0.6516 0.8994 0.5492 0.4586 0.8469
Traditional Regulation In Energy Market     0 0 0 0 0 0
5-Year Average Labor Rate     48022 48022 43282 48022 30384 41540
Social Vulnerability Index     0.3835 0.3926 0.3886 0.4156 0.4266 0.4339
@@
Number of Hazardous Facilities in 5 Miles     88 0 3 0 0 2
No Fault Lines     0 0 1 0 1 1
No Landslide Area     1 1 1 1 0 0
Peak Ground Acceleration Lower Than 0.3g     0 0 1 0 1 1
Not Having A Flood in Previous 100 Years     0 0 1 1 1 1
No Open Water Or Wetland Intersection     1 1 1 1 1 1
@@
Population Center Distance     1.5768 18.0903 22.4049 57.4029 209.4981 7.9357
Retiring Facility Distance     3.1422 39.3740 133.3787 71.0272 205.2916 0
Nuclear R&D Centers in 100 Miles     66 3 0 0 1 1
Electricity Substation Distance     1.6407 5.4444 0.5557 0.1435 0.2077 0.0277

### v2

@@ Acknowledgment
This work is sponsored by the Department of Energy Office of Nuclear Energy’s Distinguished Early Career Program (Award number: DE-NE0009424), which is administered by the Nuclear Energy University Program (NEUP). This research also made use of Idaho National Laboratory computing resources for generating preliminary results, which are supported by the Office of Nuclear Energy of the U.S. Department of Energy under Contract No. DE-AC07-05ID14517.
@@
Appendix A Site Characteristics of the Top Coal and Brownfield Sites
@@
Registry ID     C2914 C2712 C2367 C8042 110039000000 110015000000
Type     Coal Coal Coal Coal Brownfield Brownfield
Longitude     -81.4682 -79.0731 -70.7842 -80.0603 -80.32 -122.10
Latitude     40.5201 36.4833 43.0978 36.2811 27.42 37.67
County & State     Tuscarawas, OH Person, NC Rockingham, NH Stokes, NC St. Lucie, FL Alameda, CA
Siting Metric     1.0000 0.7991 0.7025 0.6650 0.6169 0.6018
State Nuclear Restrictions     0 0 0 0 0 1
State Electricity Price     11.9544 14.0500 12.6467 12.1533 13.3900 24.2231
State Net Electricity Imports     37952 14875 -6623 14875 13226 75504
State Nuclear Inclusive Policy     1 1 0 1 0 1
Population Sentiment of Nuclear Energy     0.0741 0.4586 0.0300 0.0904 0.5901 0.4019
Traditional Regulation In Energy Market     1 0 1 1 1 0
5-Year Average Labor Rate     43920 30384 38740 30384 31532 48022
Social Vulnerability Index     0.4111 0.4266 0.4347 0.4139 0.5058 0.3835
@@
Number of Hazardous Facilities in 5 Miles     2 0 5 0 2 88
No Fault Lines     1 1 1 1 1 0
No Landslide Area     1 0 1 0 1 1
Peak Ground Acceleration Lower Than 0.3g     1 1 1 1 1 0
Not Having A Flood in Previous 100 Years     1 1 0 1 0 0
No Open Water Or Wetland Intersection     0 1 1 1 1 1
@@
Population Center Distance     132.511 209.498 68.441 23.753 2.458 1.577
Retiring Facility Distance     0 205.292 78.225 0 32.546 3.142
Nuclear R&D Centers in 100 Miles     0 1 2 0 0 66
Electricity Substation Distance     0.0269 0.2077 0.1330 0.2162 0.1530 1.6407